# Polarization of reflected X-ray emission from Sgr A molecular complex: multiple flares, multiple sources?

### v1

@@
Polarization of reflected X-ray emission from
@@
Extended X-ray emission observed in the direction of several molecular clouds in the Central Molecular Zone (CMZ) of our Galaxy exhibits spectral and temporal properties consistent with the “X-ray echo” scenario. It postulates that the observed signal is a light-travel-time delayed reflection of a short (1.5 yr) and bright () flare, most probably produced a few hundred years ago by Sgr A*. This scenario also predicts a distinct polarization signature for the reflected X-ray continuum, with the polarization vector being perpendicular to the direction towards the primary source and polarization degree (PD) being determined by the scattering angle. We report the results of two deep observations of the currently brightest (in reflected emission) molecular complex Sgr A taken with the Imaging X-ray Polarimetry Explorer (IXPE) in 2022 and 2023. We confirm the previous polarization measurement for a large region encompassing Sgr A complex with higher significance, but also reveal an inconsistent polarization pattern for the brightest reflection region in its center. Specifically, X-ray polarization from this region is almost perpendicular to the expected direction in the case of Sgr A* illumination and shows a smaller degree of polarization compared to the large region. Taken at face value, this could indicate the simultaneous propagation of several illumination fronts throughout the CMZ, with the origin of one of them not being Sgr A*. The primary source could be associated with the Arches stellar cluster or a currently unknown source located in the closer vicinity of the illuminated cloud, potentially lowering the required luminosity of the primary source. Although significantly deeper observations with IXPE would be required to unequivocally distinguish between the scenarios, a combination of high-resolution imaging and micro-calorimetric spectroscopy offers an additional promising path forward.
@@ 1 Introduction
X-ray reflection off molecular clouds in the Central Molecular Zone (CMZ) of our Galaxy offers a unique window into the past activity record of the currently dormant supermassive black hole Sgr A* (Vainshtein & Syunyaev, 1980; Sunyaev et al., 1993; Markevitch et al., 1993; Koyama et al., 1996; Ponti et al., 2013). The light-travel-delay due to propagation from the primary source to the reflector implies that the currently observed reflected X-ray emission is connected to the primary’s source luminosity decades to centuries ago (as first realized and described for nova eruptions by Kapteyn, 1901; Couderc, 1939).
@@
The apparent variability of the observed diffuse emission (Revnivtsev et al., 2004; Muno et al., 2007; Inui et al., 2009; Ponti et al., 2010; Capelli et al., 2012; Clavel et al., 2013; Ryu et al., 2013; Zhang et al., 2015; Krivonos et al., 2017; Chuard et al., 2018; Terrier et al., 2018; Kuznetsova et al., 2022; Khabibullin et al., 2022; Brunker et al., 2025; Stel et al., 2025) can be best explained by the short flare scenario – the duration of the primary energetic event responsible for the clouds’ illumination cannot exceed the observed duration of reflected emission from the smallest sub-structures, the light-crossing time111The ‘scanning’ speed of the illumination front propagation along the line of sight varies from at the back side of the illumination paraboloid to infinity in its front side ( is the speed of light). of which is relatively short (e.g., Churazov et al., 2017a).
@@
Since the observed surface brightness of the reflected emission is determined primarily by the distance between the cloud and the source, the mean density of the molecular gas and flare’s fluence (flux integrated over duration of a flare), one can use density estimates obtained from molecular line emission to obtain a lower limit on the average luminosity of the primary source (corresponding to the upper limit on the flare’s duration for a fixed value of fluence). Such estimates give us a limit that still allows a flare from an X-ray binary at the Eddington level, , lasting for one year as a possible origin for the X-ray echo. On the other hand, for an Eddington level flare of Sgr A*, , the duration of merely an hour would suffice to produce the required X-ray fluence. Although large amplitude flares are regularly observed from Sgr A* in infrared and X-ray monitoring campaigns (Eckart et al., 2004; Dodds-Eden et al., 2009; Ponti et al., 2017), the fluence required to account for the X-ray echo is larger by many orders of magnitude and it is unclear if it could have been produced by the same mechanism (e.g., Ponti et al., 2013; Churazov et al., 2017c).
@@
An important and instrumental prediction of the X-ray reflection scenario is linear polarization of the continuum part of the emission, with the polarization degree determined (in the simplest case of single scattering and unpolarized primary emission) by the cosine of the scattering angle and the polarization angle (PA) being perpendicular to the direction towards the primary emission source (Churazov et al., 2002, 2017a; Marin et al., 2014, 2015; Khabibullin et al., 2020b; Di Gesu et al., 2020; Ferrazzoli et al., 2021). Thus, X-ray polarimetry has been long recognized as a potential way to pin-down the location of the primary source, the relative 3D position of the reflecting cloud, and help break degeneracies in parameters of the spectral reflection model (e.g., relative abundance of iron responsible for production of the brightest fluorescent line at 6.4 keV, Churazov et al., 2017b).
@@
The first polarimeteric observation of the currently brightest (in reflected emission) molecular complex Sgr A (Khabibullin et al., 2022; Brunker et al., 2025; Stel et al., 2025) was performed by the Imaging X-ray Polarimetry Explorer (IXPE, Weisskopf et al., 2022) observatory in 2022, broadly confirming the prediction for a short flare from Sgr A* scenario some 200 yr ago (Marin et al., 2023). Even for this brightest complex and rather long effective exposure time of the observation, the level of statistical and systematic noise enabled a polarization measurement only for the whole complex and detection of the polarization at a 3 level.
@@
In order to improve the statistical significance of this result and enable measurements for individual substructures, a second observation was performed in 2023, doubling the total clean (after filtering time intervals affected by enhanced background level) exposure time at the source position (Churazov et al., 2024).
@@
Here we report the results of the two epochs of deep observations of Sgr A taken with IXPE. We confirm the previous polarization measurement for a large region encompassing the Sgr A complex with higher significance, but also reveal a possibly inconsistent polarization pattern for the brightest reflection region in its center. Specifically, X-ray polarization from this bright region is almost perpendicular to the expected direction in the case of Sgr A* illumination and shows a smaller degree of polarization compared to the large region.
@@
The paper is organized as follows: we outline the short flare scenario and expected polarization signal in Sect. 2. The Chandra and IXPE data are described in Sect. 3. In Sect. 4, we demonstrate consistency of the observed emission from the Sgr A region with the X-ray reflection scenario using epoch-to-epoch variations of the diffuse emission. We present spatially-resolved spectro-polarimetric measurements from the combined IXPE dataset in Sect. 5. We discuss the results in Sect. 6, and present the summary and outlook in Sect. 7.
@@ 2 Short flare scenario
We use the coordinate system such that the primary source is located at the origin, the -axis is pointing from the observer towards it, and the reflecting cloud resides in the -plane (see Fig. 1). This plane is referred to as the scattering plane in what follows, and the -axis represents the direction normal to it.
The line-of-sight and image plane coordinates of the reflecting gas are connected by the relation
which can be formulated in the dimensionless (and time-independent) form after dividing by the characteristic length scale pc , so that :
@@
For the flare duration yr , the thickness of the gas layer contributing to the reflected emission for any line-of-sight direction is small
so one can neglect variations in the scattering geometry across this layer.
The degree of polarization depends on the cosine of the scattering angle (in the limit of Thomson scattering) as
Thus, knowing the degree of polarization , one can determine the 3D position of an illuminated cloud. For instance, measured for the Sgr A complex by Marin et al. 2023 corresponds to , and for one gets , i.e. the illuminated cloud is further away from us than Sgr A∗ by this distance.
One can explicitly relate the delay time and the measured degree of polarization
This expression will be used to estimate the age of the flare.
An extension of this basic scenario, allowing for possible intrinsic polarization of the original flare’s emission, was considered by Churazov et al. (2002), who presented explicit expressions for the polarization degree of the scattered emission as a function of the cloud’s relative position (see also Gilfanov et al., 1987, where scattering of anisotropic and polarized time-variable radiation has been modeled). This analysis was expanded by Khabibullin et al. (2020b) to predict the full polarization pattern of the scattered emission, i.e., the maps of the polarization degree and polarization plane orientation for various configurations of the primary source polarization. In contrast to the case of the unpolarized primary emission, the predicted patterns demonstrate substantial morphological complexity, including non-axisymmetric variations in the intensity, polarization degree, and polarization plane orientation. For instance, in this case, the electric field vector is not necessarily perpendicular to the line connecting the primary source and the reflecting cloud. Also, given that, contrary to the scattered continuum emission, fluorescent line emission is insensitive to the intrinsic polarization of the initial flare, correlated spatial variations in polarization and spectral properties (e.g., equivalent width of the fluorescent line) are expected in this case (Khabibullin et al., 2020b).
@@ 3 Data
High spatial resolution X-ray images were obtained with the Chandra Observatory (Weisskopf et al., 2000) over multiple observing campaigns from 2000-2022. The total exposure of the Chandra pointings with G0.13–0.11 located in the central region of the field of view (FOV) with high spatial resolution is 1.4 Msec. The Chandra data reduction follows a standard procedure based on the latest versions of the data reduction software (CIAO v. 4.14) and calibration (CALDB v. 4.9.8). Our particular approach and analysis steps are described in detail in Vikhlinin et al. (2009).
Briefly, they include the identification and removal of high background periods, the correction of photon energies for the time and detector temperature dependence of the charge transfer inefficiency and gain, and the creation of matching background datasets using blank sky observations with exposure times similar to the GC pointings. For the analysis presented here, we used the combined flat-fielded and background-subtracted Chandra images in the 3–8 keV band and spectra extracted in several regions of interest. Following the standard approach for analyzing Chandra spectra of extended sources, we have generated the spectral response files that combine the position-dependent ACIS calibration with the weights proportional to the observed brightness.
@@
In this work, we used the data of two IXPE observations of the Galactic Center region, more specifically, a complex of molecular clouds degrees to the east of Sgr A* (Marin et al., 2023) – the Sgr A complex. These observations with a total clean (i.e., after excluding periods of enhanced particle background and solar flares) exposure time of Msec, were performed in two parts – in February 2022 and September 2023. These two observations were targeted at (RA, Dec)=(266.51) and (266.57), respectively, with an intentional offset between their aim-points of .
@@
The raw IXPE data were processed with the standard pipeline (version 1.9.2 with HEASOFT v6.34)222https://heasarc.gsfc.nasa.gov/docs/software/heasoft/ and CALDB version of May 2025. The output FITS files of this pipeline contain the event-by-event Stokes parameters (see Kislat et al., 2015) from which the polarization observables of the X-ray radiation can be derived. The data products are publicly available for use by the international astrophysics community at the High-Energy Astrophysics Science Archive Research Center333https://heasarc.gsfc.nasa.gov/docs/ixpe/archive/ (HEASARC, at the NASA Goddard Space Flight Center). An energy-dependent particle background rejection algorithm (IXPE-Background444https://github.com/aledimarco/IXPE-background tool, v2.3) was applied to these level-2 event files; it allows the removal of of the instrumental background (Di Marco et al., 2023). Observation periods affected by increased background due to solar activity were removed (see Marin et al. 2023 for additional details on the procedure). The second epoch observation was performed with a smaller dithering amplitude, aiming at achieving a higher sensitivity in the central part of the field of view. As a result, the brightest compact (i.e., spatially unresolved) source, the Arches cluster, which was used to adjust the astrometry of the first observation, was not visible in the nominal field of view. The Arches cluster was, however, detectable near a corner of the field of view of Detector Unit 1 (DU1) in the Level 1 (L1) event data. Custom analysis of L1 data to determine attitude offsets for the individual observation segments yielded a means for consistent astrometric correction. These attitude corrections were applied after the standard pipeline processing to generate the final Level 2 (L2) products, used here and available in the HEASARC. The resulting astrometry is fully sufficient for robustly singling out the emission of the brightest reflection region, as well as pulsar wind nebula (PWN) G0.13-0.11. No additional weighting or binning of the data was included in the spatial (broad-band) or spectropolarimetric analyses. Long IXPE observations of the Circinus galaxy were used to model quiescent background spectrum, i.e. the sum of the sky and detector backgrounds in the area not contaminated by the target source (after applying identical filtering routines), and subtract it in the spectral analysis of the Stokes data (consistently with the procedures performed and described in Marin et al. 2023).

### v2

# Polarization of reflected X-ray emission from the Sgr A molecular complex: Multiple flares, multiple sources?

@@
Polarization of reflected X-ray emission from the
@@
The extended X-ray emission observed in the direction of several molecular clouds in the central molecular zone (CMZ) of our Galaxy exhibits spectral and temporal properties consistent with the X-ray echo scenario. This concept postulates that the observed signal is a light-travel-time delayed reflection of a short (1.5 yr) and bright () flare that was most probably produced a few hundred years ago by Sgr A*. This scenario also predicts a distinct polarization signature for the reflected X-ray continuum, with the polarization vector being perpendicular to the direction toward the primary source and the polarization degree being determined by the scattering angle. We report the results of two deep observations of the currently brightest (in reflected emission) molecular complex Sgr A taken with the Imaging X-ray Polarimetry Explorer (IXPE) in 2022 and 2023. We confirm the previous polarization measurement for a large region encompassing the Sgr A complex with high significance. We also reveal an inconsistent polarization pattern for the brightest reflection region in its center. Specifically, the X-ray polarization from this region is almost perpendicular to the expected direction in the case of Sgr A* illumination, and it shows a smaller degree of polarization compared to the large region. Taken at face value, this could indicate the simultaneous propagation of several illumination fronts throughout the CMZ, with the origin of one of them not being Sgr A*. The primary source could be associated with the Arches stellar cluster or a currently unknown source located closer to the illuminated cloud, potentially lowering the required luminosity of the primary source. Although significantly deeper observations with IXPE would be required to unequivocally distinguish between the scenarios, a combination of high-resolution imaging and micro-calorimetric spectroscopy offers an additional promising path forward.
@@ 1 Introduction
X-ray reflection off molecular clouds in the central molecular zone (CMZ) of our Galaxy offers a unique window into the past activity record of the currently dormant supermassive black hole Sgr A* (Vainshtein and Syunyaev, 1980; Sunyaev et al., 1993; Markevitch et al., 1993; Koyama et al., 1996; Ponti et al., 2013). The light-travel-delay due to propagation from the primary source to the reflector implies that the currently observed reflected X-ray emission is connected to the primary’s source luminosity from decades to centuries ago (as first realized and described for nova eruptions by Kapteyn, 1901; Couderc, 1939).
@@
The apparent variability of the observed diffuse emission (Revnivtsev et al., 2004; Muno et al., 2007; Inui et al., 2009; Ponti et al., 2010; Capelli et al., 2012; Clavel et al., 2013; Ryu et al., 2013; Zhang et al., 2015; Krivonos et al., 2017; Chuard et al., 2018; Terrier et al., 2018; Kuznetsova et al., 2022; Khabibullin et al., 2022; Alboslani et al., 2025; Brunker et al., 2025; Stel et al., 2025) can be best explained by the short flare scenario. It posits that the duration of the primary energetic event responsible for the clouds’ illumination cannot exceed the observed duration of reflected emission from the smallest sub-structures, the light-crossing time111The ”scanning” speed of the illumination front propagation along the line of sight varies from at the back side of the illumination paraboloid to infinity in its front side ( is the speed of light). of which is relatively short (e.g., Churazov et al., 2017a).
@@
Since the observed surface brightness of the reflected emission is determined primarily by the distance between the cloud and the source, i.e., the mean density of the molecular gas and flare’s fluence (flux integrated over duration of a flare), one can use density estimates obtained from molecular line emission to obtain a lower limit on the average luminosity of the primary source (corresponding to the upper limit on the flare’s duration for a fixed value of fluence). Such estimates give us a limit that still allows a flare from an X-ray binary at the Eddington level, , lasting for one year as a possible origin for the X-ray echo. On the other hand, for an Eddington-level flare of Sgr A*, , the duration of merely an hour would suffice to produce the required X-ray fluence. Although large amplitude (up to a factor of more than one hundred relative to the quiescent level; see Nowak et al., 2012; Haggard et al., 2019) flares are regularly observed from Sgr A* in infrared and X-ray monitoring campaigns (Eckart et al., 2004; Dodds-Eden et al., 2009; Ponti et al., 2017), the total energy required to account for the X-ray echo is larger by many orders of magnitude ( erg versus erg), and it is unclear if it could have been produced by the same mechanism (e.g., Ponti et al., 2013; Churazov et al., 2017c).
@@
An important and instrumental prediction of the X-ray reflection scenario is linear polarization of the continuum part of the emission, with the polarization degree determined (in the simplest case of single scattering and unpolarized primary emission) by the cosine of the scattering angle and the polarization angle (PA) being perpendicular to the direction toward the primary emission source (Churazov et al., 2002, 2017a; Marin et al., 2014, 2015; Khabibullin et al., 2020a; Di Gesu et al., 2020; Ferrazzoli et al., 2021). Thus, X-ray polarimetry has long been recognized as a potential way to pin down the location of the primary source and the relative 3D position of the reflecting cloud and help break degeneracies in parameters of the spectral reflection model (e.g., relative abundance of iron responsible for production of the brightest fluorescent line at 6.4 keV, Churazov et al., 2017b).
In a more complex scenario, when the X-ray reflection throughout the CMZ is produced not by a single flare but by a set of flares separated in time, or even by multiple independent sources, the spatial pattern of the X-ray reflection signal offers the cleanest way to build the complete illumination picture (e.g., Marin et al., 2014, 2015). Finally, intrinsic linear polarization of the primary flare emission might have characteristic imprints on the spatial polarization properties even in the simplest case of a single flare illumination (e.g., Khabibullin et al., 2020b). Given that different scenarios predict a variety of temporal, spectral, and polarization relations (driven by dependence of the continuum albedo on the scattering angle and independence of the fluorescent line emission of the geometrical configuration; Churazov et al., 2002), it is crucial to obtain a sensitive spectrally resolved polarization map rather than a single image-integrated quantity.
@@
The first polarimeteric observation of the currently brightest (in reflected emission) molecular complex Sgr A (Khabibullin et al., 2022; Brunker et al., 2025; Stel et al., 2025) was performed by the Imaging X-ray Polarimetry Explorer (IXPE; Weisskopf et al., 2022) observatory in 2022, and it broadly confirmed the predicted scenario of a flare from Sgr A* some 200 years ago (Marin et al., 2023). Even for this brightest complex and rather long effective exposure time of the observation, the level of statistical and systematic noise enabled a polarization measurement only for the whole complex and detection of the polarization at a 3 level.
@@
In order to improve the statistical significance of this result and enable measurements for individual substructures, a second observation was performed in 2023. This doubled the total clean exposure time (after filtering time intervals affected by enhanced background level) at the source position (Churazov et al., 2024).
@@
We report the results of the two epochs of deep observations of Sgr A taken with IXPE. We confirm the previous polarization measurement for a large region encompassing the Sgr A complex with high significance, but we also reveal a possibly inconsistent polarization pattern for the brightest reflection region in its center. Specifically, X-ray polarization from this bright region is almost perpendicular to the expected direction in the case of Sgr A* illumination, and it shows a smaller degree of polarization compared to the large region.
@@
The paper is organized as follows: We outline the short flare scenario and expected polarization signal in Sect. 2. The Chandra and IXPE data are described in Sect. 3. In Sect. 4, we demonstrate the consistency of the observed emission from the Sgr A region with the X-ray reflection scenario using epoch-to-epoch variations of the diffuse emission. We present spatially resolved spectro-polarimetric measurements from the combined IXPE dataset in Sect. 5. We discuss the results in Sect. 6, and we present the summary and outlook in Sect. 7.
@@ 2 Short flare scenario
We use the coordinate system such that the primary source is located at the origin, the -axis is pointing from the observer toward it, and the reflecting cloud resides in the -plane (see Fig. 1). This plane is referred to as the scattering plane in what follows, and the -axis represents the direction normal to it.
The line of sight and image plane coordinates of the reflecting gas are connected by the relation
which can be formulated in the dimensionless (and time-independent) form after dividing by the characteristic length scale pc so that :
@@
For the flare duration yr , the thickness of the gas layer contributing to the reflected emission for any line-of-sight direction is small,
so one can neglect variations in the scattering geometry across this layer. The degree of polarization, , depends on the cosine of the scattering angle, , (in the limit of Thomson scattering) as
Thus, knowing the degree of polarization, , one can determine the 3D position of an illuminated cloud. For instance, as measured for the Sgr A complex by Marin et al. 2023 corresponds to , and for , one gets . That is, the illuminated cloud is farther away from us than Sgr A∗ by this distance.
Using Eq. (4) and Eq. (6), one can explicitly relate the age of the flare and the measured degree of polarization as
We used this expression to estimate the age of the flare.
Given that the degree of polarization depends on the scattering angle, the finite extent of the illuminated medium along the line of sight might be important for a long flare ( from Eq. (1)). This extent can be neglected as long as is much smaller than the distance between the primary source and the scattering cloud. Given that varies between (scattering cloud behind the source with ) and (scattering cloud in front of the source with ), the reflected emission from clouds closer to us than the primary source (in the single-scattering approximation) is short-lived, while the clouds farther away will remain visible for a much longer time. Specializing for the latter case, the condition is sufficient to neglect the finite extent of the cloud. In the limit of a very long flare illuminating an infinite homogeneous medium (i.e., all scattering angles are present for a given line of sight direction), the resulting polarization is (assuming dipole phase function for continuum scattering).222This can be calculated taking into account that total scattered luminosity per unit length , while the polarized one as from Eq. (3), (4) and (6), and the corresponding integrals are easily evaluated via substitution.
An extension of this basic scenario, allowing for possible intrinsic polarization of the original flare’s emission, was considered by Churazov et al. (2002), who presented explicit expressions for the polarization degree of the scattered emission as a function of the cloud’s relative position (see also Gilfanov et al., 1987, where scattering of anisotropic and polarized time-variable radiation has been modeled). This analysis was expanded by Khabibullin et al. (2020a) to predict the full polarization pattern of the scattered emission, i.e., the maps of the polarization degree and polarization plane orientation for various configurations of the primary source polarization. In contrast to the case of the unpolarized primary emission, the predicted patterns demonstrate substantial morphological complexity, including non-axisymmetric variations in the intensity, polarization degree, and polarization plane orientation. For instance, in this case, the electric field vector is not necessarily perpendicular to the line connecting the primary source and the reflecting cloud. Also, given that, contrary to the scattered continuum emission, fluorescent line emission is insensitive to the intrinsic polarization of the initial flare, correlated spatial variations in polarization and spectral properties (e.g., equivalent width of the fluorescent line) are expected in this case (Khabibullin et al., 2020a).
@@ 3 Data
3.1 IXPE data
@@
In this work, we used the data of two IXPE observations of the Galactic Center region, more specifically, a complex of molecular clouds degrees to the east of Sgr A* (Marin et al., 2023) – the Sgr A complex centered on (RA, Dec)=(266.566), i.e., in Galactic coordinates. These observations with a total clean (i.e., after excluding periods of enhanced particle background and solar flares) exposure time of Msec, were performed in two parts – in February 2022 and September 2023. These two observations were targeted at (RA, Dec)=(266.51) and (266.57), respectively, with an intentional offset between their aim-points of .
@@
The raw IXPE data were processed with the standard pipeline (version 1.9.2 with HEASOFT v6.34)333https://heasarc.gsfc.nasa.gov/docs/software/heasoft/ and CALDB version of May 2025. The output FITS files of this pipeline contain the event-by-event Stokes parameters (see Kislat et al., 2015) from which the polarization observables of the X-ray radiation can be derived. The data products are publicly available for use by the international astrophysics community at the High-Energy Astrophysics Science Archive Research Center444https://heasarc.gsfc.nasa.gov/docs/ixpe/archive/ (HEASARC, at the NASA Goddard Space Flight Center). An energy-dependent particle background rejection algorithm (IXPE-Background555https://github.com/aledimarco/IXPE-background tool, v2.3) was applied to these level-2 event files; it allows the removal of of the instrumental background (Di Marco et al., 2023). Observation periods affected by increased background due to solar activity were removed (see Marin et al. 2023 for additional details on the procedure). The second epoch observation was performed with a smaller dithering amplitude, aiming at achieving a higher sensitivity in the central part of the field of view (FoV). As a result, the brightest compact (i.e., spatially unresolved) source, the Arches cluster, which was used to adjust the astrometry of the first observation, was not visible in the nominal FoV. The Arches cluster was, however, detectable near a corner of the FoV of Detector Unit 1 (DU1) in the Level 1 (L1) event data. Custom analysis of L1 data to determine attitude offsets for the individual observation segments yielded a means for consistent astrometric correction. These attitude corrections were applied after the standard pipeline processing to generate the final Level 2 products, used here and available in the HEASARC. The resulting astrometry is fully sufficient for singling out the emission of the brightest reflection region, as well as pulsar wind nebula (PWN) G0.13-0.11. No additional weighting or binning of the data was included in the spatial (broad-band) or spectropolarimetric analyses. Long IXPE observations of the Circinus galaxy were used to model quiescent background spectrum, i.e., the sum of the sky and detector backgrounds in the area not contaminated by the target source (after applying identical filtering routines), and subtract it in the spectral analysis of the Stokes data (consistently with the procedures performed and described in Marin et al. 2023).
3.2 Chandra data
High spatial resolution X-ray images were obtained with the Chandra Observatory (Weisskopf et al., 2000) over multiple observing campaigns from 2000 to 2024. The total exposure of the Chandra pointings having Sgr A complex located in the central region of the FoV is 1.5 Msec. The Chandra data reduction follows a standard procedure based on the latest versions of the data reduction software (CIAO v. 4.14) and calibration (CALDB v. 4.9.8). Our particular approach and analysis steps are described in detail in Vikhlinin et al. (2009). Briefly, they include the identification and removal of high background periods, the correction of photon energies for the time and detector temperature dependence of the charge transfer inefficiency and gain, and the creation of matching background datasets using blank sky observations with exposure times similar to the GC pointings. For the analysis presented here, we used the combined flat-fielded and background-subtracted Chandra images in the 3–8 keV band and spectra extracted in several regions of interest. Following the standard approach for analyzing Chandra spectra of extended sources, we have generated the spectral response files that combine the position-dependent ACIS calibration with the weights proportional to the observed brightness.